OpenStudy (anonymous):

9(3x – 16) + 15 = 6x – 24 Part A: Write the steps you will use to solve the equation, and explain each step. Part B: What value of x makes the equation true?

OpenStudy (anonymous):

Part A: Distribute the 9 to the 3x - 16. Combine Like Terms. Add 129 (the integer) to both sides. Subtract 6x (the variable) from both sides. Divide both sides by the variable coefficient (21). Get a value for x. Part B: 27x -144+15 = 6x -24 27x - 129 = 6x -24 27x = 6x+ 105 21x = 105 x = 5